# CostBench: Evaluating Multi-Turn Cost-Optimal Planning and Adaptation in Dynamic Environments for LLM Tool-Use Agents

### v1

@@ Abstract
Current evaluations of Large Language Model (LLM) agents primarily emphasize task completion, often overlooking resource efficiency and adaptability. This neglects a crucial capability: agents’ ability to devise and adjust cost-optimal plans in response to changing environments. To bridge this gap, we introduce CostBench, a scalable, cost-centric benchmark designed to evaluate agents’ economic reasoning and replanning abilities. Situated in the travel-planning domain, CostBench comprises tasks solvable via multiple sequences of atomic and composite tools with diverse, customizable costs. It also supports four types of dynamic blocking events, such as tool failures and cost changes, to simulate real-world unpredictability and necessitate agents to adapt in real time. Evaluating leading open-sourced and proprietary models on CostBench reveals a substantial gap in cost-aware planning: agents frequently fail to identify cost-optimal solutions in static settings, with even GPT-5 achieving less than 75% exact match rate on the hardest tasks, and performance further dropping by around 40% under dynamic conditions. By diagnosing these weaknesses, CostBench lays the groundwork for developing future agents that are both economically rational and robust. [^1] Our code is publicly available at https://github.com/JiayuJeff/CostBench.
@@
Jiayu Liu1   Cheng Qian2   Zhaochen Su1   Qing Zong1   Shijue Huang1   Bingxiang He3 Yi R. (May) Fung1 1Hong Kong University of Science and Technology 2University of Illinois Urbana-Champaign    3Tsinghua University {jliufv,yrfung}@ust.hk    chengq9@illinois.edu
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in reasoning DeepSeek-AI et al. (2025); Wang et al. (2025f), code generation Mai et al. (2025); Zhoubian et al. (2025), and complex problem-solving Yu et al. (2025); Dou et al. (2025). When equipped with external tools, they can interact with dynamic environments such as the web Qiao et al. (2025); He et al. (2025) and interactive systems Qin et al. (2024a); Su et al. (2025a). This integration of tool use gives LLMs agentic capabilities, enabling them to autonomously execute multi-step tasks.
@@
UserBench (Qian et al., 2025b)
MCP-Bench (Wang et al., 2025g)
@@ 2 Related Work
Figure 1: Overview of the CostBench pipeline. Starting from high-quality queries generated from combinations of user preferences, the agent constructs its plan, then interacts with an environment set up with atomic and composite tools under flexible cost assignments (atomic tool costs are randomized between 15 and 25 in our experiments), and executes actions along an customizable dynamic blocking module to achieve its goal.
@@
Recent work emphasizes evaluating LLMs under economically grounded conditions. Existing tool-use benchmarks such as MINT Wang et al. (2024b), ToolQA Zhuang et al. (2023), and MTU-Bench Wang et al. (2025d) focus on task completion but ignore efficiency. Cost-related benchmarks like PlanBench Valmeekam et al. (2023), SayCanPay Hazra et al. (2024), and ACPBench Hard Kokel et al. (2025, 2024) consider cost only in single-turn planning, limiting applicability to dynamic, iterative settings. Multi-turn benchmarks such as UserBench Qian et al. (2025b) and TravelPlanner Xie et al. (2024) focus on candidate prices rather than operational costs, while others like R-ConstraintBench Jain and Wetter (2025) and Flex-TravelPlanner Oh et al. (2025) emphasize performance–resource trade-offs under fixed budgets. Most also conflate cost with path functionality and use static environments, obscuring cost sensitivity. In contrast, CostBench decouples cost from functional bias through randomized tool assignments and equivalent solution paths, while introducing dynamic blocking events to assess adaptive planning. A detailed comparison with prior work is provided in Table 1.
@@
Prior work has examined cost-awareness in agents from several angles. Number of tool call represents the simplest form of tool cost which is thoroughly examined and optimized through algorithms (Qian et al., 2025a; Wang et al., 2025c, b). Furthermore, Toolformer Schick et al. (2023) focuses on when to invoke tools rather than reasoning about fine-grained tool costs or long-term expenditure. Starting from FrugalGPT Chen et al. (2024b), many studies further explore multi-agent coordination to reduce expenses, achieving cost control through selective and hierarchical model invocation Zhang et al. (2024); Panda et al. (2025); Gandhi et al. (2025), often in limited or static scenarios. SayCanPay Hazra et al. (2024) and CATP-LLM Wu et al. (2024) explicitly optimize for cost but remain limited to one-off planning without dynamic adaptation to changing environments. Other studies treat cost merely as token usage and focus on efficient reasoning under token constraints Han et al. (2025b); Wen et al. (2025); Huang et al. (2025b), neglecting broader operational costs. Overall, these works target narrow aspects of cost optimization, whereas our benchmark significantly broadens the notion of cost and systematically evaluates agent adaptability under realistic, dynamically changing conditions and diverse practical constraints.
@@ 3 CostBench
Moreover, CostBench can operate in two distinct modes: a static mode, where user preferences, tool costs, and tool availability remain constant; and a dynamic mode, where these attributes evolve in response to the agent’s behavior or environmental changes, capturing the uncertainty inherent in real-world scenarios. Each task unfolds through an iterative interaction loop in which the agent observes the current task state, reasons about its next action, invokes relevant tools, and receives cost-tagged feedback that updates its internal state. This continual feedback enables adaptive decision-making and dynamic replanning until the task goal is achieved. We present the detailed design of this framework in the following.
@@ 3.1 Environment Setup
In this work, we focus on the travel planning domain. Following UserBench Qian et al. (2025b) and TravelPlanner Xie et al. (2024), we categorize travel-related activities into six domains: Location, Transportation, Accommodation, Attraction, Dining, and Shopping. Each task in CostBench follows a structured four-stage workflow comprising preference identification, candidate search, candidate filtering, and final selection. Among these, the filtering stage involves multiple steps, while the other three are modeled as single-step operations. We define the task sequence as the total number of steps across all stages, providing a clear metric for task length and complexity. This multi-step design enhances the scalability of the benchmark, enabling the flexible generation of tasks with varying lengths and levels of difficulty.
Task Name
Pref. Comb.
Train Split
Filtered
Test Split
Filtered
@@
Table 2: Main statistics of query construction in CostBench. We generate 6,000 training and 1,000 test user preference combinations and filter out those violating common sense. The total number of user preference combinations equal the sum of both splits, with a discussion of the slightly imbalanced distribution provided in Appendix D.1.
Tools Value
Task Sequence NN
Total Atomic Tools NN
Total Composite Tools N(N1)/2{N*(N-1)/2}
Total Tools N(N+1)/2{N*(N+1)/2}
Table 3: Main statistics of tools used in CostBench. The variable NN denotes a adjustable task sequence, and the total numbers of atomic and composite tools vary accordingly, as indicated (we use $N=5$ in most of our experiments).
@@
To ensure interpretability and enable precise progress tracking, we introduce a set of self-defined data types that explicitly specify the input–output schema for each tool. This design prevents agents from skipping intermediate reasoning steps through direct guessing and allows the environment to automatically verify whether each action meaningfully contributes to the final solution. In CostBench, each task step maps to an atomic tool performing a non-separable operation, while a composite tool is a higher-level tool executing a sequence of atomic tools. For example, “Search_and_Decide_Location” combines “Decide_Location_Preference” and “Search_Location_Preference” into an independent tool equivalent to running them sequentially (See Figure 1). All tools follow the sequential order of task steps, with each tool consuming the outputs of preceding steps as its inputs (See example in Figure 1). This sequential dependency enables any consecutive subset of atomic tools to form a valid composite tool, thereby supporting scalability and the construction of longer or more complex tool chains while maintaining explicit input–output relationships. Details of tool construction and an example of the tool schema is illustrated in Appendix B.1 and Figure 13.
@@
Each atomic tool is assigned a randomized cost drawn from a customizable range, while the cost of a composite tool is defined as the sum of its constituent atomic tools plus Gaussian noise with zero mean and a customizable standard deviation. Tool costs are independently resampled for every data instance but remain reproducible, ensuring that environment variations are diverse yet comparable. For tasks sharing the same sequence length, cost assignments are kept consistent across static and dynamic blocking settings to ensure fair comparison. Further implementation details are provided in Appendix B.6.1.
@@ 3.2 Agent–Environment Interaction
CostBench includes an optional dynamic blocking module that introduces runtime perturbations to evaluate the robustness of cost-aware agents. Unlike static interventions, blocking events are triggered adaptively during execution in response to the agent’s behaviors. Specifically, the agent may encounter four types of runtime disruptions:
@@
The mechanism ensures that: (1) Dynamic and Uniform Triggering: Blocking events are triggered dynamically based on the estimated length of the cost-optimal path, leading to approximately uniform trigger positions and ensuring that disruptions occur at meaningful stages. (2) Independence Across Datapoints: For each run, environments are instantiated independently across datapoints, preventing correlations among blocking events. (3) Randomized Yet Reproducible Execution: Even under dynamic conditions, the experiments remain randomized yet reproducible, preserving both diversity and fairness in evaluation (see Appendix B.6).
@@ 3.3 User Query Construction
For each of the six travel-planning tasks, we construct user queries through a multi-stage pipeline. Each task defines four preference dimensions with ten candidate values, where four are used for test query construction and six for the training set, ensuring clear differentiation and non-overlapping splits, resulting in $4^{4}+6^{4}=1{,}552$ unique user preference combinations per task. All combinations are verified by GPT-4o and partially checked manually for clarity (see Appendix E). Natural language queries are generated by GPT-4o to capture realistic vagueness while preserving clear distinctions among user preferences within each dimension. An example query is shown in Figure 12, and dataset statistics are summarized in Tables 2 and 3.
Model Name Cost Path User Intent Tool Call
Cost Gap $\downarrow$ AED $\downarrow$ ANED (%) $\downarrow$ EMR (%) $\uparrow$ UIHR (%) $\uparrow$ ITUR (%) $\downarrow$
@@
Table 4: CostBench main evaluation results for task sequences of length 5 in the static setting. Scores in bold indicate the best performance among all models, and underlined scores denote the second-best performance. The metrics in (%) are in percentage forms. Reported values under Cost Gap indicate the average cost gap, with numbers in parentheses computed after excluding samples with redundant tool calls (see Appendix D.2 for details). The greedy baseline selects the atomic tool with the lowest average cost at each step (See Appendix B.2).
Figure 2: Models’ average normalized edit distance and exact match ratio for task sequences of length five to eight. Models show significantly lower performance as task complexity increases. In the most challenging case with task sequence length of eight, even the strongest model, GPT-5, achieves less than 75% exact match ratio.
@@ 4.1 Settings
We evaluate both proprietary and open-source models to ensure a balanced and comprehensive assessment of current LLM capabilities. Proprietary models include GPT OpenAI (2024b), Claude Anthropic (2025), Deepseek DeepSeek-AI et al. (2025), and Gemini Google DeepMind (2025); while open-source models include GLM Team et al. (2025), Qwen3 Qwen Team (2025), and Llama3 Touvron et al. (2023). All models use a temperature of 0.0 for deterministic decoding and a maximum token length of 16,384 to prevent output truncation.
@@
We evaluate performance using metrics across four main categories: cost, path, user-intent understanding, and tool use. The majority of these metrics quantify the deviation of an agent’s trajectory from a ground-truth trajectory, with a greedy policy serving as a baseline for comparison. Detailed definitions of all metrics are provided in Appendix B.3.
(1) Cost Gap: This metric directly measures the gap between gt accumulative costs and agent’s accumulative cost.
@@
(4) Exact Match Ratio (EMR) (%): This metric evaluates the ratio of datapoints that agent implements an identical tool call trajectory as gt.
(5) User Intent Hit Ratio (UIHR) (%): This metric reflects whether the agent correctly understands the user intent in the query. Note that there is only one correct answer for each datapoint.
(6) Invalid Tool-Use Ratio (ITUR) (%) This metric evaluates the proportion of valid tool use. Invalid behaviour includes calling wrong tool name or calling wrong parameter name, call a tool but doesn’t actually have the input type, etc.
Prompts and Environment Settings
@@ 4.2 Results
Our main results in Table 4 reveal several notable trends. First, only GPT-5 and Gemini-2.5-Pro show strong exact-match performance (95.52% and 83.73%, respectively) at a task length of five, whereas all other models fall below 60%. Performance further declines for longer task sequences (see Section 5.1), where even GPT-5 drops to below 75% EMR. This highlights a substantial gap in executing complete cost-aware plans, as even the strongest model, GPT-5, falls short of reliably producing optimal trajectories. Second, compared to the greedy baseline, some LLMs fail to show convincing improvements. For example, Qwen3-8B, GPT-4o, and Deepseek-V3.1 yield nearly identical NED scores to greedy policy, with GPT-4o even exhibiting a higher ED, indicating that its generated plans deviate further from optimal solution than greedy baseline. More concerningly, Llama-3.1-8B-Instruct underperforms the greedy baseline across all path-related metrics, suggesting fundamental weaknesses in cost-optimal planning.
@@
Comparison of Open-Source and Proprietary Models.
Beyond the greedy baseline, we compare open-source and proprietary models. Interestingly, some proprietary models show no clear advantage Chen et al. (2025): Qwen3-32B consistently surpasses GPT-4o and DeepSeek-V3.1 on all path-related metrics, and GLM-4.5 achieves performance comparable to Claude-Sonnet-4.0. While partly due to differing invalid tool-use rates, our metrics decouple tool use failure from cost/path measurements (see Appendix B.3.2), indicating that some proprietary models’ shortcomings in cost-sensitive planning are substantive and reflect genuine limitations in their reasoning capabilities.

### v2

@@ Abstract
Current evaluations of Large Language Model (LLM) agents primarily emphasize task completion, often overlooking resource efficiency and adaptability. This neglects a crucial capability: agents’ ability to devise and adjust cost-optimal plans in response to changing environments. To bridge this gap, we introduce CostBench, a scalable, cost-centric benchmark designed to evaluate agents’ economic reasoning and replanning abilities. Situated in the travel-planning domain, CostBench comprises tasks solvable via multiple sequences of atomic and composite tools with diverse, customizable costs. It also supports four types of dynamic blocking events, such as tool failures and cost changes, to simulate real-world unpredictability and require agents to adapt in real time. Evaluating leading open-sourced and proprietary models on CostBench reveals a substantial gap in cost-aware planning: agents frequently fail to identify cost-optimal solutions in static settings, with even GPT-5 achieving less than 75% exact match rate on the hardest tasks, and performance further drops significantly under dynamic conditions. By diagnosing these weaknesses, CostBench lays the groundwork for developing future agents that are both economically rational and robust.  [^1] Our code is publicly available at https://github.com/JiayuJeff/CostBench.
@@
Jiayu Liu1   Cheng Qian2   Zhaochen Su1   Qing Zong1   Shijue Huang1   Bingxiang He3 Yi R. (May) Fung1 1Hong Kong University of Science and Technology 2University of Illinois Urbana-Champaign    3Tsinghua University jliufv@connect.ust.hk    yrfung@ust.hk
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in reasoning DeepSeek-AI et al. (2025); Wang et al. (2025e), code generation Mai et al. (2025); Zhoubian et al. (2025), and complex problem-solving Yu et al. (2025); Dou et al. (2025). When equipped with external tools, they can interact with dynamic environments such as the web Qiao et al. (2025); He et al. (2025) and interactive systems Qin et al. (2024a); Su et al. (2025). This integration of tool use gives LLMs agentic capabilities, enabling them to autonomously execute multi-step tasks.
@@
UserBench (Qian et al., 2025c)
MCP-Bench (Wang et al., 2025f)
@@ 2 Related Work
Figure 1: Overview of the CostBench pipeline. Starting from high-quality queries generated from combinations of user preferences, the agent constructs its plan, then interacts with an environment set up with atomic and composite tools under flexible cost assignments (atomic tool costs are randomized between 15 and 25 in our experiments), and executes actions along a customizable dynamic blocking module to achieve its goal.
@@
Recent work emphasizes evaluating LLMs under economically grounded conditions. Existing tool-use benchmarks such as MINT Wang et al. (2024b), ToolQA Zhuang et al. (2023), and MTU-Bench Wang et al. (2025d) focus on task completion but ignore efficiency. Cost-related benchmarks like PlanBench Valmeekam et al. (2023), SayCanPay Hazra et al. (2024), and ACPBench Hard Kokel et al. (2025, 2024) consider cost only in single-turn planning, limiting applicability to dynamic, iterative settings. Multi-turn benchmarks such as UserBench Qian et al. (2025c) and TravelPlanner Xie et al. (2024) focus on candidate prices rather than operational costs, while others like R-ConstraintBench Jain and Wetter (2025) and Flex-TravelPlanner Oh et al. (2025) emphasize performance–resource trade-offs under fixed budgets. Most also conflate cost with path functionality and use static environments, obscuring cost sensitivity. In contrast, CostBench decouples cost from functional bias through randomized tool assignments and equivalent solution paths, while introducing dynamic blocking events to assess adaptive planning. A detailed comparison with prior work is provided in Table 1.
@@
Prior work has examined cost-awareness in agents from several angles. Number of tool call represents the simplest form of tool cost which is thoroughly examined and optimized through algorithms (Qian et al., 2025b; Wang et al., 2025c, b). Furthermore, Toolformer Schick et al. (2023) focuses on when to invoke tools rather than reasoning about fine-grained tool costs or long-term expenditure. Starting from FrugalGPT Chen et al. (2024b), many studies further explore multi-agent coordination to reduce expenses, achieving cost control through selective and hierarchical model invocation Zhang et al. (2024); Panda et al. (2025); Gandhi et al. (2025), often in limited or static scenarios. SayCanPay Hazra et al. (2024) and CATP-LLM Wu et al. (2024) explicitly optimize for cost but remain limited to one-off planning without dynamic adaptation to changing environments. Other studies treat cost merely as token usage and focus on efficient reasoning under token constraints Han et al. (2025b); Wen et al. (2025); Huang et al. (2025b), neglecting broader operational costs. Overall, these works target narrow aspects of cost optimization, whereas our benchmark significantly broadens the notion of cost and systematically evaluates agent adaptability under realistic, dynamically changing conditions and diverse practical constraints.
@@ 3 CostBench
Moreover, CostBench supports evaluation in two modes: a static mode, where user preferences, tool costs, and availability remain constant; and a dynamic mode, where these attributes evolve with agent behavior or environmental changes, capturing the uncertainty in real-world scenarios. Each task unfolds through an iterative interaction loop in which the agent observes the task state, reasons about its next action, invokes relevant tools, and receives cost-tagged feedback that updates its internal state. This continual feedback enables adaptive decision-making and dynamic replanning until the task goal is achieved. We describe the detailed design below and discuss the significance and generalization of CostBench in Appendix E.
@@ 3.1 Environment Setup
In this work, we focus on the travel planning domain. Following TravelPlanner Xie et al. (2024) and UserBench Qian et al. (2025c), we categorize travel-related activities into six domains: Location, Transportation, Accommodation, Attraction, Dining, and Shopping. Each task in CostBench follows a structured four-stage workflow comprising preference identification, candidate search, candidate filtering, and final selection. Among these, the filtering stage involves multiple steps, while the other three are modeled as single-step operations. We define the task sequence as the total number of steps across all stages, providing a clear metric for task length and complexity. This multi-step design enhances the scalability of the benchmark, enabling the flexible generation of tasks with varying lengths and levels of difficulty.
Task Name Pref. Comb. (Unfiltered) Train Split Filtered Test Split Filtered
@@
Table 2: Main statistics of query construction in CostBench. We generate 6,000 training and 1,000 test user preference combinations and filter out those violating common sense. The total number of user preference combinations equals the sum of both splits, with a discussion of the slightly imbalanced distribution provided in Appendix D.1.
@@
To enable precise progress tracking, we introduce a set of self-defined data types that explicitly specify and strictly enforce the input–output schema for each tool. This design prevents agents from skipping intermediate tool calling steps through direct guessing and allows the environment to automatically verify whether each action meaningfully contributes to the final solution. In CostBench, each task step maps to an atomic tool performing a non-separable operation, while a composite tool is a higher-level tool executing a sequence of atomic tools. For example, “Search_and_Decide_Location” combines “Decide_Location_Preference” and “Search_Location_Preference” into an independent tool equivalent to running them sequentially (See Figure 1). All tools follow the sequential order of task steps, with each tool consuming the outputs of preceding steps as its inputs (See example in Figure 1). This sequential dependency enables any consecutive subset of atomic tools to form a valid composite tool, thereby supporting scalability and the construction of longer or more complex tool chains while maintaining explicit input–output relationships. Details of tool construction and an example of the tool schema are illustrated in Appendix B.1 and Figure 14.
@@
Each atomic tool is assigned a randomized cost drawn from a customizable range, while the cost of a composite tool is defined as the sum of its constituent atomic tools plus Gaussian noise with zero mean and a customizable standard deviation. Tool costs are independently resampled for every data instance but remain reproducible, ensuring environment variations are diverse yet comparable. For tasks sharing the same sequence length, cost assignments are kept consistent across static and dynamic blocking settings to ensure fair comparison. Further implementation details are provided in Appendix B.6.1.
@@ 3.2 Agent–Environment Interaction
CostBench includes an optional dynamic blocking module that introduces behavior-adaptive runtime perturbations to evaluate the robustness of cost-aware agents. The agent may encounter four types of disruptions, while preserving goal reachability (see Appendix B.5):
@@
The mechanism ensures that: (1) Dynamic and Uniform Triggering: Blocking events are triggered dynamically based on estimated cost-optimal path lengths, yielding approximately uniform trigger positions and ensuring disruptions occur at meaningful stages (see Appendix B.4). (2) Evaluation Fairness: The mechanism ensures that any agent following the ground-truth optimal path encounters the same blocking events as others, guaranteeing fair comparison across models. (3) Independence Across Data Instances: For each run, environments are instantiated independently across data points, preventing correlations among blocking events. (4) Randomized Yet Reproducible Execution: Even under dynamic conditions, experiments remain randomized yet reproducible, preserving diversity and fairness in evaluation (see Appendix B.6).
@@ 3.3 User Query Construction
For each of the six travel-planning tasks, we construct user queries through a multi-stage pipeline. Each task defines four preference dimensions with ten candidate values, where four are used for test query construction and six for the training set, ensuring clear differentiation and non-overlapping splits, resulting in $4^{4}+6^{4}=1{,}552$ unique user preference combinations per task. All combinations are verified by GPT-4o and partially checked manually for clarity (see Appendix F). Natural language queries are generated by GPT-4o to capture realistic vagueness while preserving clear distinctions among user preferences within each dimension. An example query is shown in Figure 13, and dataset statistics are summarized in Tables 2 and 4.
Model Name Cost Path Task Completion Tool Call
Cost Gap $\downarrow$ AED $\downarrow$ ANED (%) $\downarrow$ EMR (%) $\uparrow$ TCR (%) $\uparrow$ ITUR (%) $\downarrow$
@@
Table 3: CostBench main evaluation results for task sequences of length 5 in the static setting (see results for task sequence of length 8 in Table 9). Scores in bold indicate the best performance among all models, and underlined scores denote the second-best performance. The metrics in (%) are in percentage forms. Reported values under Cost Gap indicate the average cost gap, with numbers in parentheses computed after excluding samples with redundant tool calls (see Appendix D.2 for details). The greedy baseline selects the atomic tool with the lowest average cost at each step (See Appendix B.2) and the average tool call number is provided in Table 6.
Figure 2: Models’ average normalized edit distance and exact match ratio for task sequences of length five to eight. Performance drops significantly as complexity increases. At length eight, even the strongest model, GPT-5 achieves less than 75% exact match (see Appendix C.1 for more results).
@@ 4.1 Settings
Models.
We evaluate both proprietary and open-source models to ensure a balanced and comprehensive assessment of current LLM capabilities. Proprietary models include GPT OpenAI (2024b), Claude Anthropic (2025), Deepseek DeepSeek-AI (2025), and Gemini Google DeepMind (2025); while open-source models include GLM Team et al. (2025), Qwen3 Qwen Team (2025), and Llama3 Touvron et al. (2023). All models use a temperature of 0.0 for deterministic decoding and a maximum token length of 16,384 to prevent output truncation.
@@
We evaluate performance using metrics spanning four categories: cost, path, task completion, and tool use. Most metrics measure the deviation between an agent’s trajectory and the ground-truth trajectory, with a greedy policy as the baseline. Full metric definitions and illustrative examples are provided in Appendix B.3.1 and Appendix B.3.3.
(1) Cost Gap: This metric directly measures the gap between the ground-truth and agent accumulated explicit tool costs.
@@
(4) Exact Match Ratio (EMR) (%): The ratio of data points that agent implements an identical tool call trajectory as ground truth.
(5) Task Completion Ratio (TCR) (%): The proportion of cases where the agent returns the unique correct answer, regardless of tool path.
(6) Invalid Tool-Use Ratio (ITUR) (%) This metric evaluates the proportion of invalid tool use. Invalid behavior includes using incorrect tool names, incorrect parameter names, and similar errors.
Prompts and Environment Settings.
@@ 4.2 Results
As shown in Table 3, while most models complete the task well, only GPT-5 and Gemini-2.5-Pro show strong exact-match performance at a task length of five, whereas all other models fall below 60%. Performance further declines for longer task sequences (see Section 5.1), where even GPT-5 drops to below 75% EMR. This highlights a substantial gap in executing complete cost-aware plans, as even the strongest model, GPT-5, falls short of reliably producing optimal trajectories. Second, compared to the greedy baseline, some LLMs fail to show convincing improvements. For example, Qwen3-8B, GPT-4o, and Deepseek-V3.1 yield nearly identical ANED scores to greedy policy, with GPT-4o even exhibiting a higher AED, indicating that its generated plans deviate further from optimal solution than greedy baseline. More concerningly, Llama-3.1-8B-Instruct underperforms the greedy baseline across all path-related metrics, suggesting fundamental weaknesses in cost-optimal planning. Robustness checks for our evaluation are detailed in Appendix C.2.